Equation for:
Parallel to 7x-3y=12 and passing through (-2,11)

Answer by stanbon(75887)

Parallel to 7x-3y=12 and passing through (-2,11)
------
Find slope of given line:
3y = 7x-12
y = (7/3)x - 4
slope = (7/3)
--------
Form: y = mx + b
Solve for "b":
11 = (7/3)(-2) + b
-------------------------
(33/3) = -14/3 + b
b = 47/3
---------
Equation:
y = (7/3)x + (47/3)
-------------------
Cheers,
Stan H.
========================